if sin x= 8/17, where 0 degree < x < 90 degree, determne the exact vaule of cos (x + 180)
Answer by stanbon(48558)

if sin x= 8/17, where 0 degree < x < 90 degree, determne the exact vaule of cos (x + 180)
----------
cos(x+180)=-cosx
If sinx = 8/17 , cos x = 15/17
So cos(x+180)=-15/17
cheers,
Stan H.